Swathmore Clothing Corporation grants its customers 30 days credit. The company uses the allowance method for its uncollectiReg 1 Reg 2 Req ЗА Reg 3B Prepare a summary journal entry to record the monthly bad debt accrual and the write-offs during thView transaction list Journal entry worksheet < 1 2 Record a summary entry to record the 2021 write-offs. Note: Enter debitsReg 1 Reg 2 Reg 3A Reg 3B Prepare the necessary year-end adjusting entry for bad debt expense. (If no entry is required for aComplete this question by entering your answers in the tabs below. Reg 1 Reg 2 Reg 3A Req 3B What is total bad debt expense fComplete this question by entering your answers in the tabs below. Req1 Req 2 Req 3A 1 Reg 3B How would accounts receivable a
Answer #1
1
Debit Credit
Bad debt expense 79500 =2650000*3%
      Allowance for uncollectible accounts 79500
Allowance for uncollectible accounts 42000
       Accounts receivable 42000
2
Debit Credit
Bad debt expense 4560
      Allowance for uncollectible accounts 4560
3a
Bad debt expense 84060 =79500+4560
3b
Balance sheet
Current assets:
Accounts receivable (net) 588940 =675000-86060
Workings:
Amount % uncollectible Amount uncollectible
0—60 days 385000 4% 15400
61—90 days 92000 12% 11040
91—120 days 52000 22% 11440
Over 120 days 146000 33% 48180
Total 675000 86060
Beginning balance, Allowance 44000
Add: Monthly bad debt accruals 79500
Deduct: Write-offs -42000
Unadjusted balance 81500
Required allowance 86060
Required year-end increase in allowance 4560